DIFFERENTIAL GEOMETRY/PDE SEMINAR

Wednesday, April 30, 2003 Padelford C-36 3:45 pm

Derivatives and Primatives

Washek F. Pfeffer (University of California, Davis)

A very general Gauss-Green theorem can be obtained by differentiating certain additive functions defined on the family of all bounded BV sets. The process is invariant with respect to local lipeomorphisms, and the resulting integration by parts formula can be used to study removable sets for PDEs in divergence form.

For more information about this seminar, visit the DG/PDE Seminar Web page (from the Math Department home page, www.math.washington.edu, follow the link Seminars, Colloquia, and Conferences).

The University of Washington is committed to providing access, equal opportunity and reasonable accommodation in its services, programs, activities, education and employment for individuals with disabilities. To request disability accommodation contact the Disability Services Office at least ten days in advance at: 206-543-6450/V, 206-543-6452/TTY, 206-685-7264 (FAX), or dso@u.washington.edu.